COHERENCE-DETECTED FTMW-IR SPECTROSCOPY OF CH_{3}OD IN THE OD STRETCH REGION.

\underline{SYLVESTRE TWAGIRAYEZU}, DAVID S. PERRY, Department of Chemistry, The University of Akron, Akron OH 44325; JUSTIN L. NEILL, MATT T. MUCKLE, and BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., P.O. Box 400319, Charlottesville, VA 22904.

Infrared spectra of the connected rotational levels of jet-cooled CH_{3}OD are recorded in the OD stretch region. The observed spectra in the range 2710 - 2740 cm^{−1} result from E-species transitions (3_{0} ← 3_{−1},2_{0} ← 3_{1},1_{0} ← 1_{1}) of CH_{3}OD. For the available rotational levels (K ′ = 0 , K ′ = 1 and K ′ = 2), the reduced torsional energies follow a pattern similar to the ground state. The torsional tunneling splitting in the OD stretch excited state is deduced to be 2.4 cm^{−1} as compared to 2.6 cm^{−1} in the ground state.